# A finite element solver for a thermodynamically consistent electrolyte model

### v1

@@ Abstract
In this study, we present a finite element solver for a thermodynamically consistent electrolyte model that accurately captures multicomponent ionic transport by incorporating key physical phenomena such as steric effects, solvation, and pressure coupling. The model is rooted in the principles of non-equilibrium thermodynamics and strictly enforces mass conservation, charge neutrality, and entropy production. It extends beyond classical frameworks like the Nernst–Planck system by employing modified partial mass balances, the electrostatic Poisson equation, and a momentum balance expressed in terms of electrostatic potential, atomic fractions, and pressure, thereby enhancing numerical stability and physical consistency. Implemented using the FEniCSx platform, the solver efficiently handles one- and two-dimensional problems with varied boundary conditions and demonstrates excellent convergence behavior and robustness. Validation against benchmark problems confirms its improved physical fidelity, particularly in regimes characterized by high ionic concentrations and strong electrochemical gradients. Simulation results reveal critical electrolyte phenomena, including electric double layer formation, rectification behavior, and the effects of solvation number, Debye length, and compressibility. The solver’s modular variational formulation facilitates its extension to complex electrochemical systems involving multiple ionic species with asymmetric valences.
@@ 1 Introduction
Over the years, several mesoscopic frameworks have been proposed to model the complex interaction between ion transport and electrochemical phenomena in electrolyte systems. One of the most widely used is the classical Nernst–Planck (NP) model, which describes ion fluxes as driven by concentration gradients and electric fields through diffusion and electromigration [37, 38, 42, 43]. Although effective for dilute electrolytes under near-equilibrium conditions, the NP model overlooks important physical effects such as finite ion sizes, ion–ion correlations, and thermodynamic coupling. These limitations can lead to nonphysical results, including negative concentrations and energy inconsistency, especially in systems with high ionic strength or strong fields [28, 9, 22]. The Poisson–Nernst–Planck (PNP) model extends the NP formulation by coupling it with the Poisson equation, enabling the electrostatic potential to evolve self-consistently with local charge distributions [26]. This framework is widely used to simulate ion transport in nanopores, membranes, and biological ion channels. However, the standard PNP model shares many of the NP model’s shortcomings and does not adequately capture steric effects, finite ion volumes, or non-ideal mixing—key factors in concentrated or strongly coupled systems [39, 17].
To overcome the limitations inherent in classical ion transport models—such as the neglect of ion–ion interactions, finite ion sizes, and lack of thermodynamic consistency—several advanced formulations have been developed. One notable class of models replaces the traditional flux–force relations with those derived from chemical potentials, resulting in volume-exclusion models that incorporate non-idealities through concentration-dependent chemical potentials [51, 30, 36, 54]. These approaches provide improved physical fidelity by accounting for steric effects and finite ion volumes. However, they also introduce analytical and computational complexities, particularly in systems near thermodynamic equilibrium or under strong coupling conditions.
@@
Several important contributions have advanced this framework. De Groot et al. [7] emphasized the role of finite ion sizes in establishing thermodynamically consistent relationships between chemical potential and concentration. Dreyer et al. [9] introduced models that rigorously couple mechanical forces with ionic diffusion, successfully addressing multiple deficiencies of the classical NP model. Roubíček [48] proposed a comprehensive coupled model integrating the Poisson equation, heat equation, NP equations, and the Navier–Stokes system. However, this formulation did not fully conform to the second law of thermodynamics, particularly when coupling the $N-1$ diffusion fluxes.
@@
To improve numerical treatment, Prohl and Schmuck [46] developed and analyzed two fully discrete, convergent schemes for the incompressible Navier–Stokes/Poisson–Nernst–Planck (NS/PNP) system. In the context of solid electrolytes, Braun et al. [6] presented a thermodynamically consistent model for ionic transport and derived a semi-analytical solution for a one-dimensional stationary case, highlighting how system parameters influence the formation of space-charge layers under applied voltage. Fuhrmann [12] extended the classical PNP framework to include modified formulations with enhanced physical realism and improved numerical properties. In a follow-up work, Fuhrmann [13] introduced a fully thermodynamically consistent model for an isothermal, incompressible ionic mixture in mechanical equilibrium, incorporating finite ion sizes and varying solvation numbers to provide a more comprehensive description of transport phenomena in electrolyte systems. Recently, Ankur et al. [3] proposed a finite element approach for a modified NS/PNP model, which incorporates a bulk modulus parameter to enable seamless transition between compressible and incompressible regimes. Their study demonstrates the applicability of the model across a range of conditions, including variations in temperature and fluid compressibility.
@@
From a computational standpoint, the literature addressing the numerical approximation of modified Poisson–Nernst–Planck (PNP) and Navier–Stokes/PNP (NS/PNP) systems with thermodynamic consistency remains relatively limited. In contrast, there is a substantial body of work dedicated to the classical PNP equations and their coupling with fluid flow through the NS equations [44, 27, 58, 47, 45, 49, 50]. Among various numerical methods, the finite element method (FEM) has proven to be a versatile and effective approach for solving these coupled systems. Several advancements have extended classical models to improve physical realism. For example, He and Sun [21] introduced a modified NS/PNP formulation in which the standard Poisson equation is replaced by a fourth-order elliptic equation to better capture electrostatic interactions. Sun et al. [52] developed a fully nonlinear Crank–Nicolson FEM scheme for the PNP equations, while Gao and He [14] proposed a linearized FEM with optimal error estimates. Despite these contributions, the strong coupling between electrochemical transport and hydrodynamics still presents analytical and numerical challenges, particularly in establishing well-posedness and ensuring stability for generalized or thermodynamically consistent formulations. The existence and uniqueness of solutions for the steady-state PNP system have been explored in early theoretical works [24, 40], with further extensions to coupled NS/PNP systems [25]. To enhance numerical robustness and flexibility, various alternative discretization strategies have been proposed, mostly relying on finite element discretizations using both conforming and discontinuous Galerkin methods [15, 19, 20, 29, 31, 32, 33, 57]. Despite notable progress in electrolyte modeling, the development of efficient, stable, and thermodynamically consistent numerical solvers for modified electrolyte systems remains a significant challenge. Classical approaches often fail to preserve key physical properties such as energy dissipation, entropy production, and charge neutrality, especially under nonlinear or strongly coupled conditions. These limitations motivate the present study, which aims to address these gaps through a robust computational framework.
@@
In this manuscript, we present a FEniCS based finite element solver for a thermodynamically consistent electrolyte model, designed to accurately capture electrochemical transport while preserving the physical structure of the governing equations. The model is derived from a variational formulation rooted in the principles of non-equilibrium thermodynamics, ensuring mass conservation, charge neutrality, and entropy consistency. By leveraging the FEM flexibility, the proposed solver is capable of handling complex geometries, heterogeneous boundary conditions, and multi-physics couplings. To demonstrate the effectiveness of the solver, we validate it against benchmark problems and compare its predictions with available analytical and experimental results. Furthermore, we explore the influence of key model parameters—including ionic diffusivities, spatial variations in dielectric permittivity, and externally applied fields—on electrolyte dynamics. These studies confirm the solver’s ability to deliver accurate, stable, and physically consistent solutions, thereby offering a powerful tool for advancing the simulation of electrochemical systems.
@@
The remainder of this paper is organized as follows. In Section 2, we describe the governing equations and their derivation based on non-equilibrium thermodynamics. Section 3 presents the details of the finite element implementation, including discretization strategies. Section 4 provides validation and benchmark results for the developed solver and model performance while Section 5 presents the results and discussion in details. Finally, Section 6 summarizes our findings and suggests potential future directions for extending this work.
@@ 2.1 Notation and preliminaries
$[\rho_{\alpha},n_{\alpha}]:\Omega\times\mathbb{R}^{+}\rightarrow\mathbb{R}^{+},\quad[\bm{v}_{\alpha},\bm{E}]:\Omega\times\mathbb{R}^{+}\rightarrow\mathbb{R}^{3}.$ (1)
@@ 2.2 Thermodynamically consistent electrolyte model
To address the limitations of classical ion transport models, a thermodynamically consistent electrolyte model is formulated based on non-equilibrium thermodynamics [9]. The governing equations respect conservation of mass, momentum, and electrostatic constraints while ensuring non-negative entropy production. The model captures essential phenomena including electrochemical coupling, mechanical stress, finite ion volume effects, and polarization.
@@
which require constitutive relations for the stress tensor $\bm{\sigma}$ and mass fluxes $\bm{J}_{\alpha}$, as well as, for the polarization field $\bm{P}$ as it occurs in the polarized part of the total charge density. In (10), there are $N-1$ partial mass balances because their sum yields the total mass balance. The constitutive theory defines the free energy of the system as
@@
$&\rho\psi_{\text{M}}=\left(K-p^{\text{ref}}\right)\left(1-{\textstyle\frac{n}{n^{\text{ref}}}}\right)+K\frac{n}{n^{\text{ref}}}\ln{\left({\textstyle\frac{n}{n^{\text{ref}}}}\right)},\\ &\rho\psi_{\text{E}}=nkT\sum_{\alpha=1}^{N}y_{\alpha}\ln(y_{\alpha}),\\ &\rho\psi_{\text{P}}=-\frac{1}{2}\varepsilon_{0}\chi|\bm{E}|^{2},\end{aligned}$
@@
Following Dreyer et al. [9], we consider $N-1$ independent diffusion fluxes given by
@@
Typically, the electric field is derived from a potential $\bm{E}=-\nabla\varphi$. After replacing $n^{e}$ with $n^{F}+n^{P}$, using (9) and (13) as well as inserting the expression for the stress tensor and polarization the complete system consists of $N-1$ independent partial mass balances, the total mass and momentum balances, and Poisson’s equation for the electrostatic potentials, and can be rewritten
@@ 2.3 Model simplification and reformulation
where $\nabla p=-n^{F}\nabla\varphi$ reflects the simplified momentum balance, indicating that the electric potential acts as a driving force for pressure in regions of nonzero space charge $n^{F}$. The term $-\epsilon_{0}(1+\chi)\Delta\varphi=n^{F}$ is the Poisson equation governing the electrostatic potential $\varphi$.
@@ 2.4 Comparison to classical Nernst–Planck
  • It treats all $N$ fluxes independently, ignoring the constraint that only $N-1$ of them are linearly independent due to mass conservation.
@@ 2.5.1 Solvation effect
Solvation refers to the microscopic electrostatic interactions between ionic species and surrounding solvent molecules, resulting in the formation of solvation shells. These shells consist of finite clusters of solvent molecules that arrange themselves around an ion due to local dipolar interactions [9]. The extent of solvation is characterized by the solvation number $\kappa_{\alpha}$, which denotes the average number of solvent molecules associated with each ion of species $\alpha$. A higher solvation number indicates a more pronounced solvation effect. By definition, the neutral solvent itself does not undergo solvation, leading to $\kappa_{N}=0$. The specific volume $V_{\alpha}$ of species $\alpha$ can be approximated based on its solvation number using the relation:
@@ 2.5.3 Dimensionless quantities
To facilitate analysis and numerical simulation, we introduce dimensionless variables by scaling the governing equations with appropriate reference quantities. The dimensionless variables are defined as follows [9]:
$n$ $\rightarrow n^{\text{ref}}n,\quad p\rightarrow p^{\text{ref}}p,\quad z_{\alpha}\rightarrow e_{0}z_{\alpha},\quad K\rightarrow p^{\text{ref}}K,\quad g_{\alpha}\rightarrow\frac{kT_{0}}{m_{\alpha}}g_{\alpha},$
$\mu_{\alpha}$ $\rightarrow\frac{kT_{0}}{m_{\alpha}}\mu_{\alpha},\quad n^{F}\rightarrow n^{\text{ref}}e_{0}n^{F},\quad\varphi\to\frac{kT_{0}}{e_{o}}\varphi,\quad x\rightarrow L^{\text{ref}}x.$
Using these dimensionless variables in (20) yields the dimensionless system:
@@
$\lambda^{2}\,\Delta\varphi$ $=-\,n^{F}.$
From this formulation, two key dimensionless parameters arise
@@
The parameter $\lambda^{2}$ corresponds to the scaled Debye length $\lambda_{D}^{2}=\lambda^{2}(L^{\text{ref}})^{2}=\frac{kT_{0}\epsilon_{0}(1+\chi)}{e_{0}^{2}n^{\text{ref}}}$, characterizing the thickness of the electric double layer near charged surfaces. Meanwhile, $a^{2}$ acts as a pressure scaling factor, rescaling the pressure field without influencing other physical quantities. Furthermore, the dimensionless diffusion flux is expressed as
@@
The specific Gibbs energy $g_{\alpha}$ takes different forms depending on the compressibility assumption (see Sec. 2.5.2):
Compressible case: $\quad g_{\alpha}=g_{\alpha}^{\text{ref}}+(\kappa_{\alpha}+1)a^{2}K\ln\left(1+\frac{1}{K}(p-1)\right),$ (31)
Incompressible case: $\quad g_{\alpha}=g_{\alpha}^{\text{ref}}+(\kappa_{\alpha}+1)a^{2}(p-1).$
In the dimensionless context, the free charge density is given by

### v2

@@ Abstract
In this study, we present a finite element solver for a thermodynamically consistent electrolyte model that accurately captures multicomponent ionic transport by incorporating key physical phenomena such as steric effects, solvation, and pressure coupling. The model is rooted in the principles of non-equilibrium thermodynamics and strictly enforces mass conservation, charge neutrality, and entropy production. It extends beyond classical frameworks like the Nernst–Planck system by employing modified partial mass balances, the electrostatic Poisson equation, and a momentum balance expressed in terms of electrostatic potential, atomic fractions, and pressure, thereby enhancing numerical stability and physical consistency. Implemented using the FEniCSx platform, the solver efficiently handles one- and two-dimensional problems with varied boundary conditions and demonstrates excellent convergence behavior and robustness. Validation against benchmark problems confirms its improved physical fidelity, particularly in regimes characterized by high ionic concentrations and strong electrochemical gradients. Simulation results reveal critical electrolyte phenomena, including electric double layer formation, rectification behavior, and the effects of solvation number, Debye length, and compressibility. The solver’s modular variational formulation facilitates its extension to complex electrochemical systems involving multiple ionic species with asymmetric valences. We publicly provide the documented and validated solver framework.
@@ 1 Introduction
Over the years, several mesoscopic frameworks have been proposed to model the complex interaction between ion transport and electrochemical phenomena in electrolyte systems. One of the most widely used is the classical Nernst–Planck (NP) model, which describes ion fluxes as driven by concentration gradients and electric fields through diffusion and electromigration [43, 44, 49, 50]. Although effective for dilute electrolytes under near-equilibrium conditions, the NP model overlooks important physical effects such as finite size effects, ion–ion correlations, and thermodynamic coupling. These limitations can lead to nonphysical results, including negative concentrations and energy inconsistency, especially in systems with high ionic strength or strong fields [30, 10, 24]. The Poisson–Nernst–Planck (PNP) model extends the NP formulation by coupling it with the Poisson equation, enabling the electrostatic potential to evolve self-consistently with local charge distributions [28]. This framework is widely used to simulate ion transport in nanopores, membranes, and biological ion channels. However, the standard PNP model shares many of the NP model’s shortcomings and does not adequately capture steric and finite size effects, or non-ideal mixing—key factors in concentrated or strongly coupled systems [45, 19]. The size-modified Poisson–Nernst–Planck (SMPNP) formulation addresses such limitations through additional entropic terms [40].
To overcome the limitations inherent in classical ion transport models—such as the neglect of ion–ion interactions, finite size effects, and lack of thermodynamic consistency—several advanced formulations have been developed. One notable class of models replaces the traditional flux–force relations with those derived from chemical potentials, resulting in volume-exclusion models that incorporate non-idealities through concentration-dependent chemical potentials [58, 32, 41, 61]. These approaches provide improved physical fidelity by accounting for steric and finite size effects. However, they also introduce analytical and computational complexities, particularly in systems near thermodynamic equilibrium or under strong coupling conditions.
@@
Several important contributions have advanced this framework. De Groot et al. [7] emphasized the role of finite size effects in establishing thermodynamically consistent relationships between chemical potential and concentration. Dreyer et al. [10] introduced models that rigorously couple mechanical forces with ionic diffusion, successfully addressing multiple deficiencies of the classical NP model. Roubíček [55] proposed a comprehensive coupled model integrating the Poisson equation, heat equation, NP equations, and the Navier–Stokes system. However, this formulation did not fully conform to the second law of thermodynamics, particularly when coupling the N1N-1 diffusion fluxes.
@@
To improve numerical treatment, Prohl and Schmuck [53] developed and analyzed two fully discrete, convergent schemes for the incompressible Navier–Stokes/Poisson–Nernst–Planck (NS/PNP) system. In the context of solid electrolytes, Braun et al. [6] presented a thermodynamically consistent model for ionic transport and derived a semi-analytical solution for a one-dimensional stationary case, highlighting how system parameters influence the formation of space-charge layers under applied voltage. Fuhrmann [13] extended the classical PNP framework to include modified formulations with enhanced physical realism and improved numerical properties. In a follow-up work, Fuhrmann [14] introduced a fully thermodynamically consistent model for an isothermal, incompressible ionic mixture in mechanical equilibrium, incorporating finite size effects and varying solvation numbers to provide a more comprehensive description of transport phenomena in electrolyte systems. Recently, Ankur et al. [3] proposed a finite element approach for a modified NS/PNP model, which incorporates a bulk modulus parameter to enable seamless transition between compressible and incompressible regimes. Their study demonstrates the applicability of the model across a range of conditions, including variations in temperature and fluid compressibility.
@@
From a computational standpoint, the literature addressing the numerical approximation of modified Poisson–Nernst–Planck (PNP) and Navier–Stokes/PNP (NS/PNP) systems with thermodynamic consistency remains relatively limited. In contrast, there is a substantial body of work dedicated to the classical PNP equations and their coupling with fluid flow through the NS equations [51, 29, 65, 54, 52, 56, 57]. Among various numerical methods, the finite element method (FEM) has proven to be a versatile and effective approach for solving these coupled systems. Several advancements have extended classical models to improve physical realism. For example, He and Sun [23] introduced a modified NS/PNP formulation in which the standard Poisson equation is replaced by a fourth-order elliptic equation to better capture electrostatic interactions. Sun et al. [59] developed a fully nonlinear Crank–Nicolson FEM scheme for the PNP equations, while Gao and He [15] proposed a linearized FEM with optimal error estimates. Despite these contributions, the strong coupling between electrochemical transport and hydrodynamics still presents analytical and numerical challenges, particularly in establishing well-posedness and ensuring stability for generalized or thermodynamically consistent formulations. The existence and uniqueness of solutions for the steady-state PNP system have been explored in early theoretical works [26, 47], with further extensions to coupled NS/PNP systems [27]. To enhance numerical robustness and flexibility, various alternative discretization strategies have been proposed, mostly relying on finite element discretizations using both conforming and discontinuous Galerkin methods [16, 21, 22, 31, 34, 35, 36, 64]. Recently, Ding and Zhou introduced a positivity-preserving, energy-stable second-order scheme for modified Poisson–Nernst–Planck equations [8]. Nevertheless, despite notable progress in electrolyte modeling, the development of efficient, stable, and thermodynamically consistent numerical solvers for modified electrolyte systems remains a significant challenge. Classical approaches often fail to preserve key physical properties such as energy dissipation, entropy production, and charge neutrality, especially under nonlinear or strongly coupled conditions. These limitations motivate the present study, which aims to address these gaps through a robust computational framework.
@@
In this manuscript, we present a FEniCS-based finite element solver for a thermodynamically consistent electrolyte model [10] that is derived via a variational formulation grounded in non-equilibrium thermodynamics. It enforces mass conservation, charge neutrality, and entropy consistency, and is validated against experimental measurements [11, 33]. By leveraging the FEM flexibility, the proposed solver is capable of handling complex geometries, heterogeneous boundary conditions, and multi-physics couplings. We validate the numerical convergence, stability, and accuracy of the solver against other computational benchmarks. Furthermore, we explore the influence of key model parameters, including ionic diffusivities, spatial variations in dielectric permittivity, and externally applied fields, on electrolyte dynamics. Together, these contributions yield a robust and versatile computational tool, which we release publicly to advance the understanding and design of electrochemical devices.
@@
The remainder of this paper is organized as follows. In Section 2, we describe the governing equations and their derivation based on non-equilibrium thermodynamics. Section 3 presents the details of the finite element implementation, including the discretization strategies used. Section 4 provides validation and benchmark results for the developed solver and model performance, while Section 5 provides a detailed presentation of further results and discussion. Finally, Section 6 summarizes our findings and suggests potential future directions for extending this work.
@@ 2.1 Notation and preliminaries
$\rho_{\alpha},n_{\alpha}]:\Omega\times\mathbb{R}^{+}\rightarrow\mathbb{R}^{+},\quad[\bm{v}_{\alpha},\bm{E}]:\Omega\times\mathbb{R}^{+}\rightarrow\mathbb{R}^{3}.$ (1)
@@ 2.2 Thermodynamically consistent electrolyte model
To address the limitations of classical ion transport models, a thermodynamically consistent electrolyte model is formulated based on non-equilibrium thermodynamics [10]. The governing equations respect conservation of mass, momentum, and electrostatic constraints while ensuring non-negative entropy production. The model represents finite-size (steric) saturation through ideal mixing and, optionally, solvation-induced specific volumes; explicit hard-core ionic diameters and nonlocal excluded-volume correlations are not included.
@@
which require constitutive relations for the stress tensor $\bm{\sigma}$ and mass fluxes $\bm{J}_{\alpha}$, as well as, for the polarization field $\bm{P}$ as it occurs in the polarized part of the total charge density. In (10), there are N1N-1 partial mass balances because their sum yields the total mass balance. The constitutive theory defines the free energy of the system as
@@
$\begin{aligned} &\rho\psi_{\text{M}}=\left(K-p^{\text{ref}}\right)\left(1-{\textstyle\frac{n}{n^{\text{ref}}}}\right)+K\frac{n}{n^{\text{ref}}}\ln{\left({\textstyle\frac{n}{n^{\text{ref}}}}\right)},\\ &\rho\psi_{\text{E}}=nkT\sum_{\alpha=1}^{N}y_{\alpha}\ln(y_{\alpha}),\\ &\rho\psi_{\text{P}}=-\frac{1}{2}\varepsilon_{0}\chi|\bm{E}|^{2},\end{aligned}$ (12)
@@
Following Dreyer et al. [10], we consider N1N-1 independent diffusion fluxes given by
@@
Typically, the electric field is derived from a potential $\bm{E}=-\nabla\varphi$. After replacing $n^{e}$ with $n^{F}+n^{P}$, using (9) and (14) as well as inserting the expression for the stress tensor and polarization the complete system consists of N1N-1 independent partial mass balances, the total mass and momentum balances, and Poisson’s equation for the electrostatic potentials, and can be rewritten
@@ 2.3 Model simplification and reformulation
where $\nabla p=-n^{F}\nabla\varphi$ reflects the simplified momentum balance, indicating that the electric potential acts as a driving force for pressure in regions of nonzero space charge $n^{F}$. The term $\epsilon_{0}(1+\chi)\Delta\varphi=n^{F}$ is the Poisson equation governing the electrostatic potential $\varphi$.
@@ 2.4 Comparison to classical Nernst–Planck
  • It treats all $N$ fluxes independently, ignoring the constraint that only N1N-1 of them are linearly independent due to mass conservation.
@@ 2.5.1 Solvation effect
Solvation refers to the microscopic electrostatic interactions between ionic species and surrounding solvent molecules, resulting in the formation of solvation shells. These shells consist of finite clusters of solvent molecules that arrange themselves around an ion due to local dipolar interactions [10]. The extent of solvation is characterized by the solvation number $\kappa_{\alpha}$, which denotes the average number of solvent molecules associated with each ion of species $\alpha$. A higher solvation number indicates a more pronounced solvation effect. The specific volume $V_{\alpha}$ of species $\alpha$ can be approximated based on its solvation number using the relation:
@@
The modification above is an intentionally minimal, first-order treatment of solvation in the spirit of [11]. It assumes that solvated ions occupy a larger partial specific volume than both the neutral solvent and the unsolvated ions, which restricts the applicability to solute-solvent mixtures containing a neutral solvent. Alternatively, more comprehensive approaches like [9, 14] derive more advanced models with an additional reformulation also for pressure and total number density. In the present work, we adopt the simpler approximation to keep the finite element formulation compact and robust, while leaving extensions for future study.
@@ 2.5.3 Dimensionless quantities
To facilitate analysis and numerical simulation, we introduce dimensionless variables by scaling the governing equations with appropriate reference quantities. We use $n^{\text{ref}}$, $p^{\text{ref}}$, $L^{\text{ref}}$, and the thermal voltage $kT_{0}/e_{0}$ as reference scales. The dimensionless variables (reusing the same symbols for brevity) are defined as follows [10]:
$n$ $\rightarrow n^{\text{ref}}n,\quad p\rightarrow p^{\text{ref}}p,\quad z_{\alpha}\rightarrow e_{0}z_{\alpha},\quad K\rightarrow p^{\text{ref}}K,\quad g_{\alpha}\rightarrow\frac{kT_{0}}{m_{\alpha}}g_{\alpha},$ (28)
$\mu_{\alpha}$ $\rightarrow\frac{kT_{0}}{m_{\alpha}}\mu_{\alpha},\quad n^{F}\rightarrow n^{\text{ref}}e_{0}n^{F},\quad\varphi\to\frac{kT_{0}}{e_{0}}\varphi,\quad x\rightarrow L^{\text{ref}}x.$
2.6 Summary of the model
To summarize, we consider a steady-state, dimensionless formulation on one- and two-dimensional domains with multiple ionic species. Using the dimensionless variables from (28) in (21) yields the final PDE system:
@@ 2.6 Summary of the model
$\lambda^{2}\,\Delta\varphi$ $=-\,n^{F},$
in which the two dimensionless parameters $\lambda^{2}$ and $a^{2}$ are defined as
@@
Here, $\lambda^{2}$ corresponds to the scaled Debye length $\lambda_{D}^{2}=\lambda^{2}(L^{\text{ref}})^{2}=kT_{0}\varepsilon_{0}(1+\chi)/(e_{0}^{2}n^{\text{ref}})$, characterizing the thickness of the electric double layer near charged surfaces, while $a^{2}$ is a pressure scaling factor, rescaling the pressure field without influencing other physical quantities. Furthermore, the dimensionless diffusion flux is expressed as
@@
The specific Gibbs energy $g_{\alpha}$ in (32) depends on the compressibility assumption (see Sec. 2.5.2):
[…]
while the dimensionless free charge density is given by